OpenStudy (anonymous):

Determine the leg of the 45-45-90 triangle with a hypotenuse length of 15 inches

OpenStudy (anonymous):



OpenStudy (mathstudent55):

Are you familiar with the Pythagorean theorem? \(a^2 + b^2 = c^2\) Here, a = b, and c = 15 \(a^2 + a^2 = 15^2 \) Solve for a.